Consider the following exponential generation function:

n%3D0,

where the D_n is the derangement numbers, and D0 = 1.

a. Find this formula in closed form

b. Prove the answer to (a) is correct
Answer #1

solution The Exponential generating function is given by G) = & Dux - 0 Here Dna derangement Number, Do=1 I a) To find closed. multiplying by an both sides n nco n = 0 A = { Dnach não n-)! + Z Dny och n=o. (n-i)! . To .. no = - w Now let ax)= en elb) by Expanding the A(X) = ex Ci++x++*+--- ) Le = (€ Enx) ( 1+2+x+x2+ ---) Coeff xn is given by oco - - - . CScanned with Cam! = (-+ - +?) - comparing ① & ② we get Dn = nb (ó-t. +5,- bit--- +40) clearly Do = 0!(6) - 1 Hence ® is the closed form of